OpenStudy (moongazer):

(2/21)+(1/6)=? show your solutions

OpenStudy (saifoo.khan):

11/42

OpenStudy (lalaly):

4/42 + 7/42 = 11/42

OpenStudy (moongazer):

i am getting 11/84

OpenStudy (saifoo.khan):

u r wrong.

OpenStudy (saifoo.khan):

sorry! :(

OpenStudy (moongazer):

how did the denominator became 42?

OpenStudy (lalaly):

21*2 and 6*7

OpenStudy (moongazer):

how?

OpenStudy (moongazer):

i can't really understand

OpenStudy (anonymous):

lol my bad the plusi sign looks like division its 11/42

OpenStudy (moongazer):

i get it know i just had a wrong multiplication. ^_^ sorry for that. :)

OpenStudy (anonymous):

\[\frac{2}{21}+\frac{1}{6}\]

OpenStudy (anonymous):

lcd is 42

OpenStudy (anonymous):

21*2=42 and 6*7=42

OpenStudy (anonymous):

\[\frac{2*2}{21*2}+\frac{1*7}{6*7}\]

OpenStudy (moongazer):

i got it already. thanks!

OpenStudy (moongazer):

i just had a wrong multplication

OpenStudy (anonymous):

oh- sorry. c u!